# GEDAN: Learning the Edit Costs for Graph Edit Distance

### v2

@@ E.2 Edit Cost Optimization Effect
In Figs.16 and17, we present the Principal Component Analysis (PCA)(Dunteman, 1989) results for the FreeSolv and BBBP datasets, respectively. PCA is applied to the 16 GED embedding values predicted by the networks GEDAN($\lambda=1$), GMSM ($\epsilon\geq 0$(Pellizzoni et al., 2024), GMSM ($\epsilon=0$(Pellizzoni et al., 2024), and GEDAN ($\lambda=0$). All models employ 5 layers to ensure that both architectures perceive the same graph depth. For GMSM (Pellizzoni et al., 2024), we use the retrieve method, which provides a similarity value used to construct the embeddings, whereas GEDAN directly outputs the GED, which we employ for PCA. For GEDAN ($\lambda=1$), we adopt the uniform fixed cost variant with costs equal to 1.
@@
In Figs.16 and17, each point corresponds to a graph, and its color indicates the associated label value. PCA clearly illustrates the effect of optimization on the predicted GED: optimized methods generally exhibit smoother gradient transitions, leading to a more accurate alignment with the target values. This additional experiment further demonstrates that the use of uniform costs is often poorly representative for many real-world scenarios.
Figure 16: PCA of the FreeSolv dataset (Mobley & Guthrie, 2014).
@@
Appendix F Inference Time Analysis
As discussed in the main paper, one of the limitations of our model lies in its higher resource usage, which makes it slower and less optimized compared to state-of-the-art methods. In this section, we provide a detailed analysis of the model’s inference time. All inference tests are conducted on a single GeForce RTX 4070 Ti GPU with a batch size of 64.
@@ Appendix F Inference Time Analysis
As shown in Fig. 18, our model is considerably slower than supervised models. This slowdown is primarily due to the overhead introduced by matrix initialization rather than the number of trainable parameters. In contrast, EUGENE (Bommakanti et al., 2024), which is also unsupervised, is significantly slower, taking several seconds to produce a single output. Its CPU-based optimization does not allow for acceleration of the code. Moreover, as illustrated in Figs. 12 and 13, we can train GEDAN for approximately 20 epochs before EUGENE (Bommakanti et al., 2024) completes prediction on the same dataset.
Table 16 reports the average number of trainable parameters for each model. Although GEDAN has fewer parameters than ERIC (Zhuo & Tan, 2022), EGSC (Qin et al., 2021), SimGNN (Bai et al., 2019a), and GREED (Ranjan et al., 2022), it is slower in practice. Notably, GraphEdX (Jain et al., 2024) achieves strong performance with few parameters. EUGENE (Bommakanti et al., 2024), being non-neural, has no trainable parameters. For all other models, we tune hyperparameters to maximize performance on the dataset, taking into account the differences in parameter counts.
Table 16: Average trainable parameters used to approximate the GED
SimGNN EGSC GREED ERIC GraphEdX GEDAN  (Bai et al., 2019a)  (Qin et al., 2021)  (Ranjan et al., 2022)  (Zhuo & Tan, 2022)  (Jain et al., 2024) ($\lambda=1$) ($\lambda=0$) 34,894 107,316 151,553 59,156 28,535 15,306 53,182

### v3

@@ E.4 Edit Cost Optimization Effect
In Figs.20 and21, we present the Principal Component Analysis (PCA)(Dunteman, 1989) results for the FreeSolv and BBBP datasets, respectively. PCA is applied to the 16 GED embedding values predicted by the networks GEDAN($\lambda=1$), GMSM ($\epsilon\geq 0$(Pellizzoni et al., 2024), GMSM ($\epsilon=0$(Pellizzoni et al., 2024), and GEDAN ($\lambda=0$). All models employ 5 layers to ensure that both architectures perceive the same graph depth. For GMSM (Pellizzoni et al., 2024), we use the retrieve method, which provides a similarity value used to construct the embeddings, whereas GEDAN directly outputs the GED, which we employ for PCA. For GEDAN ($\lambda=1$), we adopt the uniform fixed cost variant with costs equal to 1.
@@
In Figs.20 and21, each point corresponds to a graph, and its color indicates the associated label value. PCA clearly illustrates the effect of optimization on the predicted GED: optimized methods generally exhibit smoother gradient transitions, leading to a more accurate alignment with the target values. This additional experiment further demonstrates that the use of uniform costs is often poorly representative for many real-world scenarios.
Figure 20: PCA of the FreeSolv dataset (Mobley and Guthrie, 2014).
@@
Appendix F Computational Efficiency Analysis
As discussed in the main paper, one limitation of our model is its higher computational cost, which results in slower performance compared to state-of-the-art methods. In this section, we analyze the computational efficiency of our approach relative to existing methods, reporting both training and inference times as well as peak memory usage for each model, as summarized in Table 17.
Table 17: Computational efficiency analysis performed on a GeForce RTX 4070 Ti GPU with a batch size of 64 using FP32 precision.
Time (ms) Peak Memory (MB) Models Training Inference Training Inference Hardware SimGNN 6.9 5.9 18 10 GPU EGSC 5.2 2.6 35 25 GPU GREED 6.7 1.9 68 62 GPU ERIC 17.8 1.5 24 11 GPU GraphEdX-Xor 33.2 29.4 146 39 GPU GEDAN ($\lambda=1$) 106.3 67.6 397 134 GPU GEDAN ($\lambda=0$) 69.5 64.3 469 156 GPU EUGENE - 9222 - 512 CPU
All models except EUGENE are trained on GPU, while EUGENE runs on CPU; accordingly, its reported memory usage excludes the process-initialization overhead. All methods are evaluated using FP32 precision with a batch size of 64, ensuring a fair comparison. Models that rely on a matrix-based approximation of GED – specifically GEDAN (both supervised and unsupervised variants) and GraphEdX-Xor – exhibit higher training times and peak memory usage due to the need to allocate and maintain these matrices. Among unsupervised methods, GEDAN is generally less memory-intensive than EUGENE, whereas in the supervised setting it is substantially more memory-demanding compared to standard supervised baselines.
Table 18 reports the average number of trainable parameters for each model. Although GEDAN has fewer parameters than ERIC (Zhuo and Tan, 2022), EGSC (Qin et al., 2021), SimGNN (Bai et al., 2019a), and GREED (Ranjan et al., 2022), it is slower in practice. Notably, GraphEdX (Jain et al., 2024) achieves strong performance with few parameters. EUGENE (Bommakanti et al., 2024), being non-neural, has no trainable parameters. For all other models, we tune hyperparameters to maximize performance on the dataset, taking into account the differences in parameter counts.
Table 18: Average trainable parameters used to approximate the GED
SimGNN EGSC GREED ERIC GraphEdX GEDAN  (Bai et al., 2019a)  (Qin et al., 2021)  (Ranjan et al., 2022)  (Zhuo and Tan, 2022)  (Jain et al., 2024) ($\lambda=1$) ($\lambda=0$) 34,894 107,316 151,553 59,156 28,535 15,306 53,182
As shown in Fig. 22, our model is considerably slower than supervised models. This slowdown is primarily due to the overhead introduced by matrix initialization rather than the number of trainable parameters. In contrast, EUGENE (Bommakanti et al., 2024), which is also unsupervised, is significantly slower, taking several seconds to produce a single output. Its CPU-based optimization does not allow for acceleration of the code. Moreover, as illustrated in Figs. 13 and 14, we can train GEDAN for approximately 20 epochs before EUGENE (Bommakanti et al., 2024) completes prediction on the same dataset.
@@ Appendix F Computational Efficiency Analysis
In Fig. 23, we show the average peak memory usage observed during model inference. All models except EUGENE are trained on GPU, while EUGENE runs on CPU; accordingly, its reported memory usage excludes the process-initialization overhead. All methods are trained in FP32 precision with a batch size of 64, ensuring a fair comparison across settings.
Figure 23: Average peak memory usage during training for all models with FP32 precision and batch size 64.
Models that rely on a matrix-based approximation of GED – specifically GEDAN (both supervised and unsupervised variants) and GraphEdX-Xor – show higher memory consumption because they need to allocate and maintain these matrices. Among unsupervised methods, GEDAN is on average less memory-demanding than EUGENE, whereas in the supervised setting it is significantly more expensive in terms of memory compared to standard supervised baselines.